# Effect of Synthetic Jets Actuator Parameters on Deep Reinforcement Learning-Based Flow Control Performance in a Square Cylinder

### v2

@@ II.1 Flow environment
In the current study, the incompressible flow dynamics are analyzed using the open-source CFD toolkit, OpenFOAM, as outlined by Jasak et al Jasak et al. (2007); Jasak (2009). OpenFOAM stands as a rigorously tested and extensively applied computational framework that offers a comprehensive suite of numerical algorithms for the resolution of the Navier-Stokes equations. The solver implemented within OpenFOAM employs the finite volume method for the discretization of the computational domain, subdividing it into a structured mesh of control volumes. The numerical simulations were conducted using the pimpleFoam solver to solve the unsteady Navier-Stokes equations. The temporal discretization employed the Backward scheme, while the divergence discretization of the velocity field was implemented using the Gauss linear scheme, and the gradient discretization of the velocity utilized the leastSquares method. These discretization schemes were selected to ensure the requisite mathematical precision and stability for the simulation of fluid flow dynamics. To promote numerical stability within the simulations, the time step size is meticulously selected as $\Delta t=0.0005$.
To verify the accuracy of the grid, this study conducted a grid convergence investigation. We compared results from three different grid scales and contrasted them with computational outcomes from other studies, summarizing the findings in Table. 2. The table lists the values for the mean drag coefficient ($C_{D,Mean}$), the standard deviation of the drag coefficient ($C_{D,Std}$), the standard deviation of the lift coefficient ($C_{L,Std}$), and the $S_{t}$ at $Re=100$ and $Re=500$. All three grids utilized structured meshes, differing only in the level of refinement, with the number of grid cells detailed in the table. The computational results in Table. 2 demonstrate that the differences among the three grid resolutions are insignificant and consistent with values obtained in existing literature. In the subsequent advanced DRL training phase, the main grid scheme is employed to balance computational accuracy and cost-effectiveness, ensuring the required precision is achieved while minimizing computational overhead.
@@
Sen, Mittal, and BiswasSen, Mittal, and Biswas (2011) - 1.53 - - 0.145
Sharma and EswaranSharma and Eswaran (2004) - 1.49 - - 0.150
Singh et al.Singh et al. (2009) - 1.51 - - 0.147
Coarse 16,400 2.070 0.481 0.036 1.130
$Re=500$ Main 23,264 2.064 0.448 -0.042 0.127
Fine 28,450 2.067 0.419 -0.001 0.128
Cao, Ge, and TamuraCao, Ge, and Tamura (2012) - 2.050 - - 0.133
Bai and AlamBai and Alam (2018b) - 1.940 - - 0.136
@@ II.2 Machine learning
The agent interacts with the environment, selecting actions according to its current policy $\pi$. The experience tuples $(s_{i},a_{i},r_{i},s_{i+1})$ collected from these interactions are stored in a replay buffer. When updating the network parameters, the algorithm samples batches of experience data from the replay buffer. Moreover, entropy ($\mathcal{H}$) is a metric that quantifies the uncertainty of a random variable. Within the framework of reinforcement learning, the entropy of a policy, $\mathcal{H}(\pi(\cdot\mid s_{i}))$, represents the stochasticity of the policy at a given state $s_{i}$.Haarnoja et al. (2018a, b, 2019a) The expression for this is given by:
@@
where $\mathcal{H}(\cdot)$ is the entropy measure and the variable $\alpha$ functions as a temperature parameter that calibrates the relative significance of the entropy term against the reward, thereby modulating the stochasticity of the resulting optimal policy.Haarnoja et al. (2018b, 2019b)
@@ II.3 DRL-Enhanced Active Flow Control
The design of the DRL components is centered around this goal. The AFC problem is integrated with the DRL framework by defining three key components: the agent’s action $a_{t}$ (the mass flow rate of the synthetic jets), the agent’s state $s_{t}$ (observational data from the environment), and the agent’s reward $r_{t}$ (the control objectives). Firstly, the agent’s action, defined as the mass flow rates of the synthetic jet, is limited to a mass flow rate not exceeding 2% of the upstream inflow. Furthermore, to ensure the smoothness of continuous actions, a smoothing function $S$ is utilized between consecutive actions $a_{t}$ and $a_{t+1}$, facilitating a gradual transition and mitigating abrupt changes in jet velocity. The smoothing function $S$ is defined as follows:
@@
where $C_{D,0}$ represents the baseline drag coefficient $C_{D}$, serving as a reference point. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective being to minimize this value relative to the baseline. $(C_{L})_{T_{i}}$ denotes the lift coefficient at time step $T_{i}$. The parameter $\omega$ is a weight factor for the lift coefficient, typically ranging between 0.1 and 0.2.
@@ III.1 Baseline flow simulation
The state in which no flow control is applied is defined as the Baseline. In this study, active flow control begins from the baseline state, which also serves as a comparison to the controlled flow. As shown in Fig. 7 (a), at $Re=100$, flow separation occurs near the leading edge corners of the square cylinder. Fig. 7 (b) provides a zoomed-in view of the flow around the square cylinder, where flow separation is observed near the leading edge corners, and recirculation bubbles form behind the cylinder. In this region, the fluid flows in the reverse direction, creating closed-loop circulation. In the wake region behind the cylinder, alternating vortex shedding occurs as the recirculation bubbles oscillate. When $Re=500$, the vortex shedding pattern in the wake region of the square cylinder, as shown in Fig. 7 (c), differs from that at $Re=100$. At this higher Reynolds number, the vortices behind the cylinder are no longer alternating and regular. In Fig. 7 (d), it can be observed that flow separation occurs near the leading edge corners, creating recirculation zones on both sides of the cylinder. Behind the cylinder, a fluctuating recirculation region forms, displaying oscillatory motion in the spanwise direction.
Figure 7: Velocity magnitude of the flow around the square cylinder. (a) $Re=100$, entire domain; (b) $Re=100$, local domain; (c) $Re=500$, entire domain; (d) $Re=500$, local domain.
At $Re=100$, the wake region of the square cylinder exhibits regular vortex shedding, as shown in Fig. 8. Vortices shed from the trailing edge of the cylinder form a regular pattern, characterized by alternating vortices along the centerline, creating a street-like structure. At $Re=500$, the shedding pattern in the wake changes, as depicted in Fig. 9. After flow separation occurs at the leading edge of the square cylinder, some separated fluid reattaches to the cylinder surface. The vortex shedding in the wake region of the square cylinder becomes increasingly unstable, with vortices squeezing, stretching, and deforming as they develop downstream in the form of vortex pairs. Compared to $Re=100$, the flow instability at $Re=500$ intensifies, leading to a more complex shedding pattern in the wake. We have described the flow structures and vortex shedding patterns of the baseline flow at $Re=100$ and $Re=500$. At Reynolds number 500, the flow exhibits greater instability and complexity, indicating that the difficulty of flow control is increased. Implementing flow control techniques in such conditions requires advanced strategies that can effectively manipulate the flow dynamics and mitigate flow separation.
Figure 8: Snapshots of the spanwise vorticity field of the square cylinder wake at $Re=100$.
Figure 9: Snapshots of the spanwise vorticity field of the square cylinder wake at $Re=500$.
@@
Next, we will present the base flow simulation of a semi-cylindrical body. The base flow represents the ideal state without vortex shedding and is used solely to calculate the drag coefficient in this idealized condition. This serves as an asymptote to measure the maximum possible drag reduction. The base flow corresponds to a stable equilibrium state of the controlled Navier-Stokes equations. In this state, the velocity and pressure fields of the fluid remain unchanged, constituting a stable solution of the Navier-Stokes equations. This state can be considered a potential fixed point in the configuration space of fluid flow. However, despite the apparent stability of the baseline flow, it is, in fact, unstable to infinitesimal perturbations, leading to the occurrence of vortex shedding phenomena. Vortex shedding is one of the sources of resistance experienced by the cylinder, and our active flow control aims to mitigate this resistance. Therefore, assessing the effectiveness of the control strategy can be done by using a modified baseline resistance value, assuming the absence of vortex shedding. To simulate the scenario without vortex shedding, the strategy involves simulating only the upper half of the computational domain along the centerline. Symmetric boundary conditions are applied at the lower boundary of the computational domain (where $y=0$). These symmetric boundary conditions are explicitly defined as follows: the velocity component $v=0$, and the derivatives of the velocity $u$ and the pressure $p$ with respect to $y$ are zero, i.e., $\frac{\partial u}{\partial y}=0$ and $\frac{\partial p}{\partial y}=0$. The boundary conditions for the rest of the domain remain unchanged.This approach effectively models an idealized flow scenario.
Figure 10: The base flow is obtained through a simulation of a half-domain, where the x-coordinate ranges from -10D to 30D, and the y-coordinate ranges from 0 to 10D. We apply symmetric boundary conditions on the y = 0 boundary, and the computational domain’s mesh is consistent with Fig. 3. This configuration results in a scenario without vortex shedding, serving as a hypothetical baseline for comparison with the results of active flow control. (a) $Re=100$; (b) $Re=500$.
@@
In Fig. 10, simulations were performed for the cases of $Re=100$ and $Re=500$, where the wake flow behind the semi-square cylinder was in a steady state without vortex shedding, and the only reason for pressure drag was flow separation. The total drag coefficient encapsulates the intrinsic drag associated with the baseline flow and the correctable component, with only the second part of drag being adjustable through strategic manipulation of the von Kármán vortex sheddingProtas and Wesfreid (2002). We computed the asymptotic drag coefficients, denoted as $C_{Dh}$, for the semi-square cylinder at $Re=100$ and $Re=500$, which were found to be 0.668 and 0.498, respectively. These values were extended to obtain the baseline drag coefficients, denoted as $C_{Db}$, for a fully square cylinder in a steady-state flow, resulting in values of 1.335 and 0.995 for $Re=100$ and $Re=500$, respectively. The baseline drag coefficient, $C_{Db}$, will be compared with the drag obtained during active flow control to evaluate the effectiveness of the control strategy in reducing drag.
@@ III.2 DRL-based control performance
In the Sec. III.1, we have observed that the vortex shedding patterns are completely different at $Re=100$ and $Re=500$. This section primarily evaluates the performance of DRL-based AFC control at $Re=100$ and $Re=500$, focusing on its effectiveness in suppressing vortex shedding in the wake flow field, as well as its performance in drag reduction and lift suppression. The synthetic jets are positioned near the trailing edge corners of the square cylinder, with a width of $D/25$. In this section, we will evaluate the control performance of the DRL-based AFC technique at $Re=100$ and $Re=500$ separately. To visually observe the effect of vortex shedding suppression during the training process, Fig. 11 ($Re=100$) and Fig. 12 ($Re=500$) capture snapshots of the instantaneous velocity magnitude at four representative episodes on the learning curve. Then, we summarize the control performance in terms of drag reduction and lift suppression in Table. 3. We also compare our trained results with other research studies for comparative analysis. By examining the control performance metrics and velocity magnitude snapshots, we can gain insights into the effectiveness of the DRL-based AFC technique in suppressing vortex shedding, reducing drag, and suppressing lift at $Re=100$ and $Re=500$.
Figure 11: When $Re=100$, observed the evolution of the instantaneous velocity magnitude when DRL training. Specifically, focused on four key moments: (a) Episode = 0, (b) Episode = 900, (c) Episode = 2800, and (d) Episode = 5000.
For $Re=100$, AFC is applied to the flow around a square cylinder using DRL, and the learning curve along with snapshots of the instantaneous velocity magnitude of the square cylinder are depicted in Fig. 11. The DRL training commences with the baseline flow field, characterized by alternating vortex shedding in the wake region. Flow control is then executed based on this initial condition. The cumulative reward function curve of the DRL training exhibits a sharp increase during the initial stages. After 900 episodes of training, the cumulative reward becomes positive, indicating that the PPO agent has discovered a control strategy that meets the objective function. At this point, the vortices generated in the wake region of the baseline flow field begin to move downstream, and the previously oscillating recirculation bubble behind the square cylinder is elongated. Although the recirculation bubble continues to oscillate, no new vortices are shed. As training progresses to 2800 episodes, the growth of the cumulative reward begins to level off, indicating that the DRL training is converging. By this stage, the flow field around the square cylinder has expelled the vortices that had shed in the baseline flow, with only the initial shed vortex continuing to move towards the outlet. After 5000 episodes of training, the cumulative reward reaches a plateau, signifying that the DRL training has stabilized and fully converged. At this point, vortex shedding in the wake region of the square cylinder is completely suppressed.
@@
Similarly, under a Reynolds number of 500, we employed DRL to execute active flow control around a square cylinder, with Fig. 12 depicting its learning curve and the instantaneous velocity magnitude during characteristic episodes. In the RL training process, the cumulative reward value of the first episode ($R_{0}$) is considered the starting point of the learning process. It is observed that at $Re=500$, the $R_{0}$ of the DRL training is approximately -210, compared to -45 at $Re=100$. Additionally, the initial stage’s learning curve exhibits a steeper slope compared to the case of $Re=100$. The significant difference indicates that at higher Reynolds numbers, the initial training phase faces more severe challenges and requires more extensive exploration to discover effective control strategies. In the first episode, the shedding of vortices in the wake of the baseline flow no longer follows a regular alternating pattern but evolves jointly as "vortex pairs". After 500 episodes of training, the cumulative reward has increased to a positive value, indicating that the PPO agent has explored effective control strategies that guide the reward function towards positivity. Currently, the recirculation area in the wake of the cylinder has enlarged and developed into an approximately symmetric and stable state. The two sets of vortices originally shed by the baseline flow gradually develop downstream in the computational domain, and at this point, no new vortices are shedding in the wake of the cylinder. As DRL training progresses to 2200 episodes, the growth rate of the learning curve slows down towards convergence. At this stage, the two sets of vortices shed by the baseline flow have been mostly expelled from the computational domain, with only some vortices still being shed from the outlet of the domain. Upon reaching 4200 episodes of training, the learning curve enters a plateau phase, and the reward function tends to converge. Furthermore, the recirculation area in the wake of the cylinder has stabilized to its maximum extent, and the phenomenon of vortex shedding has been completely suppressed.
III.3 Sensitivity analyses of synthetic jets

### v3

@@ II.1 Flow environment
In the current study, the incompressible flow dynamics are analyzed using the open-source CFD toolkit, OpenFOAM, as outlined by Jasak et al. jasakLibrary; jasak2009 OpenFOAM stands as a rigorously tested and extensively applied computational framework that offers a comprehensive suite of numerical algorithms for the resolution of the Navier-Stokes equations. The solver implemented within OpenFOAM employs the finite volume method for the discretization of the computational domain, subdividing it into a structured mesh of control volumes. The numerical simulations use the pimpleFoam solver to solve the unsteady Navier-Stokes equations. The temporal discretization employs the Backward scheme, while the divergence discretization of the velocity field uses the Gauss linear scheme, and the gradient discretization of the velocity utilizes the leastSquares method. These discretization schemes are selected to ensure the requisite mathematical precision and stability for the simulation of fluid flow dynamics. To promote numerical stability within the simulations, the time step size is meticulously selected as $\Delta t=0.0005$.
To verify the accuracy of the grid, this study conducted a grid convergence investigation. We compared results from three different grid scales and contrasted them with computational outcomes from other studies, summarizing the findings in Tab. 2. The table lists the values for the mean drag coefficient ($C_{D,Mean}$), the standard deviation of the drag coefficient ($C_{D,Std}$), the standard deviation of the lift coefficient ($C_{L,Std}$), and the $S_{t}$ at $Re=100$ and $Re=500$. All three grids utilized structured meshes, differing only in the level of refinement, with the number of grid cells detailed in the table. The computational results in Tab. 2 demonstrate that the differences among the three grid resolutions are insignificant and consistent with values obtained in existing literature. In the subsequent advanced DRL training phase, the main grid scheme is employed to balance computational accuracy and cost-effectiveness, ensuring the required precision is achieved while minimizing computational overhead.
@@
sen2011flowsen2011flow - 1.530 - - 0.145
SharmaSharma - 1.490 - - 0.150
SinghSingh - 1.510 - - 0.147
Coarse 16,400 2.070 0.481 1.164 1.130
$Re=500$ Main 23,264 2.060 0.441 1.173 0.127
Fine 28,450 2.067 0.419 1.181 0.128
ShearShear - 2.050 - - 0.133
DependenceDependence - 1.940 - - 0.136
@@ II.2 Machine learning
The agent interacts with the environment, selecting actions according to its current policy $\pi$. The experience tuples $(s_{i},a_{i},r_{i},s_{i+1})$ collected from these interactions are stored in a replay buffer. When updating the network parameters, the algorithm samples batches of experience data from the replay buffer. Moreover, entropy ($\mathcal{H}$) is a metric that quantifies the uncertainty of a random variable. Within the framework of reinforcement learning, the entropy of a policy, $\mathcal{H}(\pi(\cdot\mid s_{i}))$, represents the stochasticity of the policy at a given state $s_{i}$.haarnoja2018soft; Haarnoja et al. (2018, 2019a) The expression for this is given by:
@@
where $\mathcal{H}(\cdot)$ is the entropy measure and the variable $\alpha$ functions as a temperature parameter that calibrates the relative significance of the entropy term against the reward, thereby modulating the stochasticity of the resulting optimal policy.Haarnoja et al. (2018, 2019b)
@@ II.3 DRL-Enhanced Active Flow Control
The design of the DRL components is centered around this goal. The AFC problem is integrated with the DRL framework by defining three key components: the agent’s action $a_{t}$ (the mass flow rate of the synthetic jets), the agent’s state $s_{t}$ (observational data from the environment), and the agent’s reward $r_{t}$ (the control objectives). Firstly, the agent’s actions are represented by the mass flow rates of the synthetic jets. The jet flow is controlled by dimensionless mass flow rates $Q_{1}$ and $Q_{2}$, with the condition $Q_{1}+Q_{2}=0$ to maintain zero net mass flow. This condition ensures that drag reduction is achieved through indirect flow control rather than direct momentum injection, avoiding the instability in numerical results caused by mass addition or subtraction. Additionally, the synthetic jet flow rate is limited to not exceeding 2% of the upstream inflow. Furthermore, to ensure the smoothness of continuous actions, a smoothing function $S$ is utilized between consecutive actions $a_{t}$ and $a_{t+1}$, facilitating a gradual transition and mitigating abrupt changes in jet velocity. The smoothing function $S$ is defined as follows:
@@
where $C_{D,0}$ represents the baseline drag coefficient $C_{D}$, serving as a reference point. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective being to minimize this value relative to the baseline. $(C_{L})_{T_{i}}$ denotes the lift coefficient at time step $T_{i}$. The parameter $\omega$ is a weight factor for the lift coefficient. We adopt a penalty factor of $0.2$ based on rabault2019artificialrabault2019artificial, balancing drag reduction and lift suppression.
@@ III.1 Baseline flow simulation
We define the flow state without implemented flow control as the baseline. As the flow a square cylinder exhibits many intriguing phenomena depending on the Reynolds number, we present two studies at $Re$ of 100 and 500 to investigate the mechanism of flow control, with a focus on the phenomena of vortex shedding and flow separation. Active flow control commences from the baseline state, hence, it also serves as a basis for comparison with the controlled flow. As illustrated in Fig. 7(a), at a $Re$ of 100, the flow separates right at the leading edge corner of the square cylinder. However, the separation bubble is almost negligible as the flow reattaches immediately behind the corners. The boundary layer remains largely attached to the side surface of the square cylinder, as depicted in Fig. 7(b). The primary flow separation occurs at the trailing edge of the cylinder, resulting in a wake formation predominantly at the leading (Fig. 7(c)). The wake region of the square cylinder exhibits regular vortex shedding, as shown in Fig. 7(d). Vortices shed from the trailing edge of the cylinder form a regular pattern, characterized by alternating vortices along the centerline, creating a street-like structure.
Figure 7: The streamlines and instantaneous vorticity around a square cylinder at $Re=100$. (a) Streamlines near the leading corner of the square cylinder. (b) Streamlines of the square cylinder. (c) Streamlines near the trailing corner of the square cylinder. (d) Snapshots of the spanwise vorticity field. SP represents the separation point.
In contrast, at a $Re$ of 500, the primary flow separation points shift to the leading edge corners of the cylinder. Comparing Fig. 7(a) and Fig. 8(a), flow separation occurs at the leading edge in both cases. However, there is a significant separation bubble behind the corner. As seen in Fig. 8(b) and Fig. 8(c), the separation bubble extends beyond the length of the square cylinder and engulfs the trailing edge corner, submerging it in the recirculation region. This shift suggests that at higher Reynolds numbers, the boundary layer detaches earlier due to increased inertial forces, resulting in a more turbulent wake and potentially different drag characteristics. Understanding these variations in flow separation mechanisms with changing Reynolds numbers is essential for designing effective flow control strategies for drag reduction and enhanced flow stability. At $Re=500$, the shedding pattern in the wake changes, as depicted in Fig. 8(d). After flow separation occurs at the leading edge of the square cylinder, some separated fluid reattaches to the cylinder surface. The vortex shedding in the wake region of the square cylinder becomes increasingly unstable, with vortices squeezing, stretching, and deforming as they develop downstream in the form of vortex pairs. Compared to $Re=100$, the flow exhibits greater instability and complexity at $Re=500$ , indicating an increased difficulty in flow control. In this state, implementing AFC based on DRL requires a more sophisticated control algorithm.
Figure 8: The streamlines and instantaneous vorticity around a square cylinder at $Re=500$. (a) Streamlines near the leading corner of the square cylinder. (b) Streamlines of the square cylinder. (c) Streamlines near the trailing corner of the square cylinder. (d) Snapshots of the spanwise vorticity field. SP represents the separation point, and SD represents the saddle point.
@@
To achieve our objective of suppressing unsteady vortex shedding for drag reduction, we incorporate an additional test case that simulates the conditions of a steady flow. This is achieved by simulating a half-body with symmetric boundary conditions to enforce symmetry of the wake flow. The base flow represents the ideal state without vortex shedding and is used solely to calculate the drag coefficient in this idealized condition. This serves as an asymptote to measure the maximum possible drag reduction. The base flow corresponds to a stable equilibrium state of the controlled Navier-Stokes equations where the velocity and pressure fields of the fluid remain unchanged with respect to time, constituting a steady solution of the Navier-Stokes equations. This state can be considered a potential fixed point in the configuration space of fluid flow. However, despite the apparent stability of the baseline flow, it is, in fact, unstable to infinitesimal perturbations, leading to the occurrence of vortex shedding phenomena. Vortex shedding is one of the sources of drag experienced by the cylinder, and our active flow control aims to mitigate this resistance. Therefore, assessing the effectiveness of the control strategy can be done by using a modified baseline drag value, assuming the absence of vortex shedding. To simulate the scenario without vortex shedding, the strategy involves simulating only the upper half of the computational domain along the centerline. Symmetric boundary conditions are applied at the lower boundary of the computational domain (where $y=0$). These symmetric boundary conditions are explicitly defined as follows: the velocity component $v=0$, and the derivatives of the velocity $u$ and the pressure $p$ with respect to $y$ are zero, i.e., $\frac{\partial u}{\partial y}=0$ and $\frac{\partial p}{\partial y}=0$. The boundary conditions for the rest of the domain remain unchanged. This approach effectively models an idealized flow scenario.
@@
In Fig. 9, simulations are performed for the cases of $Re=100$ and $Re=500$, where the wake flow behind the semi-square cylinder is in a steady state without vortex shedding, and the only reason for pressure drag is flow separation. The total drag coefficient encapsulates the intrinsic drag associated with the baseline flow and the correctable component, with only the second part of drag being adjustable through strategic manipulation of the von Kármán vortex sheddingprotas2002drag. We compute the asymptotic drag coefficients, denoted as $C_{Dh}$, for the semi-square cylinder at $Re=100$ and $Re=500$, which are found to be 0.668 and 0.498, respectively. These values are extended to obtain the baseline drag coefficients, denoted as $C_{Db}$, for a fully square cylinder in a steady-state flow, resulting in values of 1.335 and 0.995 for $Re=100$ and $Re=500$, respectively. The result of $Re=100$ is consistent with the study of Xia2024Xia2024. The baseline drag coefficient, $C_{Db}$, is compared with the drag obtained during active flow control to evaluate the effectiveness of the control strategy in reducing drag.
[…]
@@ III.2 DRL-based control performance
In Sec. III.1, we have observed that the vortex shedding patterns are completely different at $Re=100$ and $Re=500$. This section primarily evaluates the performance of DRL-based AFC control, focusing on its effectiveness in suppressing vortex shedding in the wake flow field, as well as its performance in drag reduction and lift suppression. The synthetic jets are positioned near the trailing edge corners of the square cylinder, with a width of $D/25$. In this section, we will evaluate the control performance of the DRL-based AFC technique at $Re=100$ and $Re=500$ separately. To visually observe the effect of vortex shedding suppression during the training process, Fig. 10 ($Re=100$) and Fig. 11 ($Re=500$) capture snapshots of the instantaneous velocity magnitude at four representative episodes on the learning curve.
Figure 10: When $Re=100$, observed the evolution of the instantaneous velocity magnitude when DRL training. Specifically, focused on four key moments: (a) Episode = 0, (b) Episode = 1000, (c) Episode = 2800, and (d) Episode = 5000.
For $Re=100$, AFC is applied to the flow around a square cylinder using DRL, and the learning curve along with snapshots of the instantaneous velocity magnitude of the square cylinder are depicted in Fig. 10. The DRL training commences with the baseline flow field, characterized by alternating vortex shedding in the wake region. Flow control is then executed based on this initial condition. The cumulative reward function curve of the DRL training exhibits a sharp increase during the initial stages. After 1000 episodes of training, the cumulative reward becomes positive, indicating that the SAC agent has discovered a control strategy that meets the objective function. At this point, the onset of vortex shedding is shifted downstream in the wake region, and the previously oscillating recirculation bubble behind the square cylinder is elongated. As training progresses to 2800 episodes, the growth of the cumulative reward begins to level off, indicating that the DRL training is converging. By this stage, the vortex shedding phenomenon has been largely suppressed, with virtually no visible vortices being shed downstream. The recirculation zone has been greatly elongated and remains relatively stable, with only the initial shed vortex continuing to move towards the outlet. After 5000 episodes of training, the cumulative reward reaches a plateau, signifying that the DRL training has stabilized and fully converged. At this point, vortex shedding in the wake region of the square cylinder is completely suppressed.
@@
Similarly, under a Reynolds number of 500, we employ DRL to execute active flow control around a square cylinder, with Fig. 11 depicting its learning curve and the instantaneous velocity magnitude during characteristic episodes. In the RL training process, the cumulative reward value of the first episode is considered the starting point of the learning process. The initial episode reward value is -300, which is much lower compared to the case with a Reynolds number of 100. This indicates that the training task at $Re=500$ is significantly more challenging. At the beginning of the control, the shedding of vortices in the wake of the baseline flow no longer follows a regular alternating pattern but evolves in the form of more complex “vortex pairs.” After 500 episodes of training, the cumulative reward increases to a positive value, indicating that the SAC agent has explored effective control strategies that guide the reward function towards positivity. Currently, the recirculation area in the wake of the cylinder has enlarged and developed into an approximately symmetric and stable state. The two sets of vortices originally shed by the baseline flow gradually develop downstream in the computational domain. As DRL training progresses to 2200 episodes, the growth rate of the learning curve slows down towards convergence. Upon reaching 4200 episodes of training, the learning curve enters a plateau phase, and the reward function tends to converge. Furthermore, the recirculation area in the wake of the cylinder has stabilized to its maximum extent, and the phenomenon of vortex shedding has been completely suppressed. To sum up, at $Re=100$ and $Re=500$, we completely suppress vortex shedding in the wake, demonstrating that the control strategy can significantly eliminate flow field instability. The control performance of DRL-based AFC is validated in the flow around a square cylinder at $Re=100$ and $Re=500$.
III.3 Parameterization and Sensitivity Analysis of Synthetic Jets